OpenStudy (crisulcampo):

factoring and express this into tg(alpha) terms: tg(alpha)sec^2 (alpha) + ctg(alpha)csc^2 (alpha) thanks

zepdrix (zepdrix):

tg? ctg?

OpenStudy (crisulcampo):

tg(alpha)

zepdrix (zepdrix):

what? that doesn't answer my question... tg(alpha) = tan(alpha) or something?

OpenStudy (agent0smith):

I assume tg is tangent and ctg is cotangent... so why not use tan and cot like a normal person :P

OpenStudy (crisulcampo):

I don't live in USA. Most of spanish math books about trygonometry express this functions like that tan (alpha) = tg(alpha) It's the same with its opposite function...

OpenStudy (crisulcampo):

I need help because I got an answer, however it doesn't match with my math book result ... So I just want to check your procedures, if you can of course... thanks

OpenStudy (agent0smith):

\[\large \tan \alpha \sec^2 \alpha + \cot \alpha \csc^2 \alpha\]use some identities \\[\large \tan \alpha (1 +\tan^2 \alpha) \alpha + \frac{ 1 }{ \tan \alpha } \left( 1+ \frac{ 1 }{ \tan^2 \alpha } \right)\]from there just simplify it

OpenStudy (mathstudent55):

You need to use identities.

OpenStudy (crisulcampo):

\[[1 + \tan^2(\alpha)][\tan^2(\alpha) + \frac{ 1 }{ \tan^3(\alpha) }]\] this is the math book answer, however I don't get this... Could you get this?

OpenStudy (agent0smith):

What an odd way of writing it :/\[ \large \tan \alpha (1 +\tan^2 \alpha) + \frac{ 1 }{ \tan \alpha } \left( 1+ \frac{ 1 }{ \tan^2 \alpha } \right)\] \[ \large \tan \alpha +\tan^3 \alpha+ \frac{ 1 }{ \tan \alpha } + \frac{ 1 }{ \tan^3 \alpha } \]

OpenStudy (crisulcampo):

It's OK, I got that you have written but, math book result is the last expression I put before... My question is Could you get the expression I put before? I mean \[[1 + \tan^2(\alpha)] [\tan^2(\alpha) + \frac{ 1 }{ \tan^3(x) }] \] Could you get this exactly?

OpenStudy (crisulcampo):

starting from your last expression \[\tan(\alpha) + \tan^3(\alpha) + \frac{ 1 }{ \tan(\alpha) } + \frac{ 1 }{ \tan^3(\alpha) }\] Could you get the one I put before? @agent0smith

OpenStudy (agent0smith):

\[\large [1 + \tan^2(\alpha)] [\tan^2(\alpha) + \frac{ 1 }{ \tan^3 \alpha }] = \] \[\large \tan^2 \alpha +\tan^4 \alpha +\frac{ 1 }{ \tan \alpha }+\frac{ 1 }{ \tan^3 \alpha }\]does not appear equivalent to what I have above; first two terms are multiplied by tan alpha.

OpenStudy (crisulcampo):

So, the math book result isn't correct! :) I was confused about my procedure but I just realized I was right! thanks